OpenStudy (anonymous):

1-sin^2x/ (sinx-cscx) a.-cos x b. -sin x c. sin2x d. cos2x

OpenStudy (adi3):

wht do u think

OpenStudy (rational):

write \(\csc x\) as \(\frac{1}{\sin x}\) and try to simplify \[\dfrac{1-\sin^2x}{\sin x-\csc x} = \dfrac{1-\sin^2x}{\sin x-\frac{1}{\sin x}} = \dfrac{\sin x(1-\sin^2 x)}{\sin^2 x-1} = -\sin x\]

OpenStudy (anonymous):

Oh I got Cos^2x...